### v1

# Vulnerabilities of smart contracts and mitigation schemes: A Comprehensive Survey

Abstract
Ethereum smart contracts are highly powerful; they are immutable and retain massive amounts of tokens. However, smart contracts keep attracting attackers to benefit from smart contract flaws and Ethereum’s unexpected behaviour. Thus, methodologies and tools have been proposed to help implementing secure smart contracts and to evaluate the security of smart contracts already deployed. Most related surveys focus on tools without discussing the logic behind them; in addition, they assess the tools based on papers rather than testing the tools and collecting community feedback. Other surveys lack guidelines on how to use tools specific to smart contract functionalities.
This paper presents a literature review combined with an experimental report, that aims to assist developers in developing secure smarts, with a novel emphasis on the challenges and vulnerabilities introduced by NFT fractionalization by addressing the unique risks of dividing NFT ownership into tradeable units called fractions. It provides a list of frequent vulnerabilities and corresponding mitigation solutions. In addition, it evaluates the community’s most widely used tools by executing and testing them on sample smart contracts. Finally, a complete guidance on how to secure smart contracts is presented.
Index Terms:
Blockchain, Ethereum, Smart contracts, Formal verification, Semantic verification, Fuzzing, Software security, Software quality.=-15pt
I Introduction
Blockchain technology has exploded in popularity over the years due to its immutability, security, and transparency in a permissionless and decentralized environment. One of the well-known and used implementations is Ethereum. As of 31 May 2023, the Ethereum cryptocurrency’s market capital surpassed $220 billion, with millions of transactions being executed everyday [1]. Ethereum uses Turing Complete blockchain technology [2], allowing developers to implement smart contracts. Smart contracts are executable programs stored in the blockchain written mostly in solidity [3]. Using smart contracts facilitates the execution of pre-defined terms without consulting third parties in an anonymous, transparent and tempered residence way.
The Ethereum blockchain is the foundation for numerous financial projects. Examples include: (a) platforms for decentralised finance (DeFi), which enable users to access a variety of financial services like lending and borrowing, trading, and insurance without the need for intermediates. (b) Stablecoins, cryptocrencies whose value is tied to that of a fiat currency or other asset, like gold or the US dollar, (c) Security token offerings (STOs), which are virtual assets that simulate ownership of a physical asset like a stock or piece of real estate. (d) Non-fungible tokens (NFTs), which are unique digital assets that might signify ownership of a physical or digital object, like a work of art or a collectible, (e) Amidst this variety, NFT fractionalization emerges as a noteworthy trend, dividing ownership of NFTs into smaller, more accessible units, thus broadening participation in the digital asset market. This value attracts attackers to exploit different vulnerabilities related to the implementation of smart contracts in order to steal cryptocurrencies or tamper with assets. For instance, a recent attack was on May 2023 where Level Finance Exchange announced the loss of more than 214,000 $LVL tokens, approximated to be $1.01 million [4], caused by an attack on its smart contract by manipulating a recursive calling vulnerability. There is a variety of vulnerabilities and causes such as arithmetic overflows, reentrancy, inadequate randomness, calling an unknown third-party contract code and failing to check the return status of external calls. The runtime environment that supports contract code execution provides more attack points. One example is when malevolent miners pick and choose which transactions are included in a mined block, or in what order they are included.
Since vulnerabilities in smart contracts can result in significant financial loss, multiple tools are proposed to identify vulnerabilities in Solidity smart-contracts. Detection methods include: (a) Symbolic execution when the program is abstractly executed to cover multiple possible inputs. (b) Fuzz testing by injecting several invalid inputs. (c)Taint analysis by checking an input flow. (d) Formal verification when the behavior of the smart contract is checked mathematically using a formal model.
In this paper we conduct a literature review combined with an experimental report. We present the most common vulnerabilities in Solidity smart contracts, including those posed by the fractionalization of NFTs and the corresponding mitigation schemes,. We also overview and compare the most popular schemes and tools used to detect vulnerabilities of smart contracts. Finally, we propose a set of guidelines for auditing smart contracts.
In this survey, we searched 2 popular databases Engineering Village and Scopus, in addition to web articles written by the Ethereum community Concerning vulnerabilities in Solidity smart-contracts, mitigation tools and guidelines. Moreover, we experimentally studied 5 tools widely used to detect vulnerabilities in smart contracts namely Oyente [5], Slither [6], Mythril [7], Manticore [8] and Echidna [9].
The rest of this paper is organised as follows: Section II outlines the methodology we used to produce this survey. Section III presents the most common vulnerabilities of smart contracts, and the corresponding mitigation schemes. Section IV presents common methodologies used to detect smart contract vulnerabilities. Section V introduces the most useful tools to detect vulnerabilities in smart contracts. Section VI presents a set of guidelines for auditing smart contracts. Section VII presents an overview of existing related surveys. Finally, Section VIII concludes the paper.
II Survey Methodology
In this section, we present the primary research questions that we hope to answer through this survey. The methodology we used to collect existing related work is then presented.
This survey aims to answer the following research questions:
RQ1: What are the most frequent vulnerabilities in Solidity smart contracts?
RQ2: How to mitigate vulnerabilities in Solidity smart contracts?
RQ3: What are the existing methodologies used to detect the vulnerabilities of smart contracts and how do they compare with each other?
RQ4: what are the most used tools by the Ethereum community to investigate/mitigate smart contracts based on Github forks and web articles?
To address these questions, we conducted a Multivocal Literature Review (MLR)[10] for data preparation; This technique incorporates both gray literature and white literature. Gray literature includes blogs, videos, and forums; it is usually written by practitioners from industry and academia. It’s not peer reviewed. White literature includes peer-reviewed research articles from journals. We carried out this step by choosing the published studies in journals, conferences with high impact factor and competitive acceptance rate. We also check the citation count of the studies being chosen on Google Scholar to evaluate their impact on the evolution of this emerging paradigm. The rationale behind choosing MLR is that the field of smart contracts security is still relatively new; thus, including literature from practitioners is helpful to have a complete overview. Moreover, the feedback provided by smart contract developers is crucial in selecting the most useful tools.
As for the search strategy, we followed a protocol presented by Kitchenham for systemic reviews [11]. This protocol consists of three phases: (a) Define a search string; (b) Use the string in search engines; and (c) Select literature based on predefined inclusion and exclusion criteria.
We defined a search string presented in Listing 1 to collect the following resources: (a) Review articles about smart contract security; (b) Papers that discuss vulnerabilities in Ethereum smart contracts; (c) Paper that that cover detection methodologies or mitigation schemes of smart contract vulnerabilities; and (d) Papers about tools that detect vulnerabilities in smart contacts. We have only included paper that contains the keywords in either the subject, title or abstract. The term ”opcode” refers to the basic instructions carried out by the blockchain’s virtual machine, which regulate the logic and operations of the smart contract. Turing completeness, on the other hand, refers to the capacity of a blockchain or smart contract language to simulate any Turing machine, which means it can resolve any computational problem given enough time and resources. Both those terms are used to represent smart contract.
Figure 2 illustrates the complete process of articles identifications. 504 conference and journal articles were found using the specified search string. Out of these, 41 articles are included based on the exclusion and inclusion criteria mentioned in table I. As for gray literature we focused on blog posts found within the first eight pages of Google search results. The search strings used included ”ethereum smart contract vulnerabilities and mitigation” and ”ethereum smart contract analysis tools”. Out of numerous blogs reviewed, eight posts were found to meet our established criteria. To assist the quality of the gray literature we have taken into consideration the following aspects:
The reputation of the publisher.
The expertise of the author in the field of smart contract security such as job position.
The content is clearly stated with supported arguments.
Criteria | Description |
Inclusion Criteria |
- Studies related to vulnerabilities in Ethereum: Vulnerabilities explication and/or mitigation techniques.
- Studies related to the security improvement of Ethereum smart contracts. - Studies related to vulnerability detection tools in Ethereum smart contracts. - Studies that introduce open-source tools to detect the vulnerabilities of solidity smart-contracts. |
Exclusion Criteria |
- Non-English papers.
- Results past the first eight Google pages, as we have noticed that after the eight pages the results are not relevant. - Data-sets, tweets, presentations. - Tools that use machine learning, as comparing machine learning tools need different criteria such as the used training models, accuracy, etc. . - Studies that are based on the improvement of blockchain infrastructure, rather than the smart contract programming - Study that is an older version of another paper that is previously been considered. |
III Vulnerabilities
In this section, we present some of the common vulnerabilities of Ethereum smart contracts alongside real world attacks and prevention mechanisms.
III-A Reentrancy
III-A1 Description
When implementing smart contracts, reentrancy is one of the most important vulnerability to address, also known as recursive call attack. A contract calling another contract, with external calls, will cause the stoppage of the calling contract’s execution and memory state until the called function returns a response. External calls are exposed in contract interfaces; hackers can use them to invoke a function within the contract numerous times, causing the contract to perform unanticipated tasks. This vulnerability occurs in a solidity smart contract performing critical tasks (e.g.,token transfer) before resolving the effects that should have been addressed (e.g., balance update).
III-A2 Implications on NFT Fractionalization
Reentrancy vulnerabilities have already manifested significant impacts on decentralized finance (DeFi) protocols, illustrating potential risks for fractionalized NFT platforms. For instance, the dForce DeFi Protocol Hack [12] in February 2023, where an attacker exploited a reentrancy vulnerability in the Curve Finance vault on the Arbitrum and Optimism blockchains, part of the dForce protocol, led to the theft of approximately $3.6 million in assets. This attack underscores the danger reentrancy poses to smart contracts handling complex financial transactions, serving as a warning for platforms dealing with fractionalized NFTs. Furthermore, the CREAM Finance Hack [13] in August 2021 and the Siren Protocol Hack [14] in September 2021, with losses of $18.8 million and $3.5 million respectively, further exemplify the ongoing threat of reentrancy attacks to the blockchain and DeFi sectors.
Reentrancy attacks can distort market dynamics by unfairly redistributing assets, leading to market manipulation and loss of liquidity. For fractionalized NFTs, where valuation depends on both the perceived value of the underlying asset and the platform’s integrity, such attacks can result in volatile price swings and diminished market confidence.
III-A3 Protection Measures
It is recommended to use send() and transfer() methods instead of the general call() method to transfer money. They are deemed safer because they have a gas limit of 2,300 gas. Whereas, the method call() does not have a gas limit and forwards the remaining gas to the target address. The gas limit prohibits the target contract from making costly external function calls. The checks-effects-interactions pattern [15] is the most reliable approach to prevent reentrancy attacks. This pattern defines how the code of a function should be organized to minimize undesirable side effects and execution behaviors. The programmer needs to include all checks which usually consist of assert() and need() modifiers. If these checks pass, the function will then resolve all of the effects of the contract ( e.g. balance update). The function should only communicate with other contracts once all state changes have been updated. Even if an attacker performs a recursive call to the initial function, the user will not be able to exploit the state of the contract since external functions are called last.
Another protection mechanism against reentrancy attacks, is the use of mutex. A mutex locks down the contract state. Mutex allows only one execution of a critical section of code at a time using lock mechanism. Only the lock’s owner has the capacity to modify it. When an attacker tries to perform a reentrancy attack, the state will be locked by the previous call and the balance can not be updated. Mutex must be treated with care to ensure that the lock can be released. ReentrancyGuard is OpenZeppelin’s own mutex implementation [16]. This library offers a nonReentrant modifier that secures an external function with a mutex.
III-A4 Example
Listing 3 illustrates a simplified smart contract, FractionalNFT, designed to manage fractional shares of an NFT’s revenue. This contract allows owners of fractional shares to withdraw their proportion of sales revenue when the NFT is sold. The contract contains a critical flaw in its withdrawShare function, which makes it susceptible to reentrancy attacks. In this contract, the withdrawShare function fails to adhere to the checks-effects-interactions pattern, updating the shareholder’s balance after transferring funds. This ordering allows for the potential reentrancy, where a malicious actor could recursively call withdrawShare within a fallback function, draining the contract’s funds beyond their rightful share.
To address this vulnerability, the ReentrancyGuard utility from the OpenZeppelin security library can be integrated to prevent recursive calls, as demonstrated in Listing 4. By employing the nonReentrant modifier provided by ReentrancyGuard, the revised contract ensures that the withdrawShare function cannot be re-entered while it is still executing, effectively mitigating the reentrancy vulnerability.
III-B Front Running
III-B1 Description
Front-running, also known as transaction ordering dependency, occurs when the execution logic depends on the order of the transactions submitted to it. The order of transactions in Ethereum is determined by the miner [17]. The transaction is visible to the network before being executed. The participants can exploit from this visibility by sending transactions with a higher gas price to be included first.
III-B2 Implications on NFT Fractionalization
Front-running can have profound implications in the context of NFT fractionalization that might include:
1. Manipulation of Share Prices: Malicious actors could exploit transaction ordering to manipulate the market for fractional tokens, buying up tokens at lower prices before a large transaction increases their value, or conversely, selling tokens to depress prices ahead of a significant sell order.
2. Interference with Auction Mechanisms: Many NFT platforms use auction mechanisms for selling fractional shares or entire NFTs. Front-runners could preemptively place bids to disrupt fair auction outcomes, affecting the final sale price of an NFT.
3. Unfair Distribution of Revenue: In scenarios where revenue from NFT sales is distributed among share owners, front-runners could strategically insert transactions to claim a disproportionate share of the distributions, undermining the fairness of the process.
III-B3 Protection Measures
The best solution to protect against front running vulnerability, is to remove the advantage of transaction ordering from the application. A solution is to remove the importance of time. Another possible solution is to use a commit-reveal hash scheme where the participant submits the hash of the answer instead of the answer. The contract then stores the hash and the address of the sender; the answer is revealed only after all all the responses are submitted.
III-B4 Example
Listing 5 illustrates a smart contract managing the sale of fractional tokens of an NFT, vulnerable to front-running. In this scenario, an attacker could observe pending purchase transactions and execute their own purchase with a higher gas fee, securing shares at the current price before a large purchase increases their value, thereby disadvantaging legitimate buyers.
To mitigate the vulnerability and ensure fair transaction processing, the contract can be revised to include a commit-reveal scheme. This modification requires buyers to commit to a purchase without revealing the exact quantity or price initially, followed by a reveal phase where the transaction details are disclosed. Mitigating the front-running attack as showing in Listing 6.
III-C Arithmetic
III-C1 Description
Solidity’s integer types, such as uint8, uint16, and uint256, are constrained by fixed sizes, capable of representing numbers within specific ranges. Arithmetic operations that result in values outside these permissible ranges lead to integer overflow (exceeding the maximum value) or underflow (dropping below zero) [18] . These vulnerabilities can cause smart contracts to behave unpredictably, potentially leading to significant security breaches or financial losses.
III-C2 Implications on NFT Fractionalization
In the context of NFT fractionalization, arithmetic vulnerabilities pose unique challenges. For instance when distributing sales revenue among shareholders, overflow or underflow errors can lead to incorrect allocation of funds. This not only affects the fairness and accuracy of fees distribution but also exposes the platform to potential exploitation. Arithmetic vulnerabilities, particularly overflows and underflows, have had substantial consequences in the decentralized finance (DeFi) sector, highlighting the associated risks for platforms dealing with fractionalized NFTs. A prominent example includes the mintToken function in the Coinstar (CSTR) Ethereum token’s smart contract [19], which suffered from an integer overflow. This flaw permitted the contract owner to arbitrarily adjust the balance of any user to any chosen value. An other example is, the 4chan gang group experienced a substantial loss of $2.3 million due to an underflow vulnerability in the ERC-20 token implementation of PoWH (Proof of Weak Hands), which allowed an attacker to exploit this flaw for financial gain [20].
III-C3 Protection Measures
In order to avoid integer overflows and underflows, arithmetic operations should be properly implemented by checking the operators and operands before performing the operation. It is recommended to use the assert() and require() modifiers. It is also advisable to use the library for arithmetic functions, called SafeMath by OpenZeppelin [18]. Solidity version 0.8.0 has included this library; if an overflow/underflow occurs, the transaction will revert.A prevalent method for detecting this type of vulnerability involves the use of taint analysis, a technique that we will elaborate on in Section IV.
III-C4 Example
Listing 7 showcases a smart contract managing fractional tokens of an NFT, vulnerable to underflow. If an attempt is made to transfer more tokens than available, under Solidity versions prior to 0.8.0, this could lead to an underflow, setting the token balance to an incorrect high value, potentially enabling malicious tokens distribution.
Given Solidity 0.8.0’s automatic checks for arithmetic operations, the primary vulnerability of underflows and overflows is directly addressed by the compiler. However, we can implement additional logic to further safeguard the integrity of fractional tokens calculations and transfers. An improved version of the smart contract is presented in Listing 8.
III-D Mishandled Exceptions
III-D1 Description
Solidity provides two primary paradigms for interacting with external contracts: direct contract calls and low-level calls. Direct contract calls involve invoking functions directly on known contract interfaces, which inherently revert on failure, thereby throwing an exception. Conversely, low-level calls, executed via methods like call(), delegatecall(), and callcode(), return a boolean success flag instead of reverting on exceptions [21]. These calls do not inherently revert transaction execution upon failure; instead, they return false, necessitating explicit checks of their return values to ensure intended execution flow. Failure to adequately check the result of a low-level call can lead to unintended execution continuation, potentially compromising contract logic and security. This vulnerability was notably exploited in the King of Ether game, where the smart contract’s failure to verify the result of a send() operation led to discrepancies in payments, resulting in users overpaying or underpaying [22].
III-D2 Implications on NFT Fractionalization
Mishandled exceptions, notably in operations involving the transfer of tokens or the distribution of revenues, pose significant risks to the platform’s functionality. For example, if a smart contract designed to distribute sales revenue from an NFT among its token holders neglects to confirm the success of these transactions, it could lead to financial disparities. Such scenarios may arise when the contract employs low-level calls for fund transfers without verifying their execution success. Malicious entities might seize on these vulnerabilities, intentionally causing transactions to fail silently by making a contract reject transactions in its fallback function.
III-D3 Example
The example 9 shows a contract that manages fractional ownership of an NFT. It includes a function to distribute proceeds from NFT sales to token holders. The contract incorrectly handles a low-level call when sending proceeds, posing a risk if the call fails.
In the revised contract 10, we use a safer approach to distribute proceeds by checking the success of each payment and reverting the transaction if a payment fails.
III-E Code Injection via delegatecall
III-E1 Description
There is a unique type of method called delegate call. The DELEGATECALL opcode is similar to a conventional message call with the exception that the code performed at the target address is executed in the context of the calling contract. The current address, storage, and balance refer to the calling contract. This enables a smart contract to load code from another smart contract dynamically at runtime. Calling into untrusted contracts via delegatecall() is particularly risky since the code at the target smart contract has complete control over the caller’s balance; thus, it can modify any of the caller’s storage data.
III-E2 Implications on NFT Fractionalization
Malicious code executed through delegatecall might tamper with the contract’s ownership or control mechanisms, enabling attackers to reroute assets or funds. Furthermore, attackers could change the revenue distribution logic, redirecting profits meant for rightful owners to unauthorized entities. In extreme cases, like the second Parity multi-sig attack where an attacker took over three main Parity wallets and stole $31 million [23] , these vulnerabilities could let attackers completely take over the NFT fractionalization contract. This means they could push out the real owners and managers, potentially locking, freezing, or stealing assets.
III-E3 Protection Measures
When employing delegatecall(), caution is paramount, especially when dealing with contracts not fully trusted. It’s critical to avoid making delegatecall() to addresses derived from user inputs unless there’s a rigorous verification process against a list of trusted contracts. This precaution helps in ensuring that only known, secure contracts can be interacted with, significantly reducing the risk of malicious interference.
Solidity’s library keyword facilitates the creation of library contracts, designed to be stateless and immune to being destroyed. By defining a contract as a stateless library, it’s guaranteed that the executing code cannot alter the storage data of the calling contract. This design principle is crucial for minimizing risks associated with storage context issues that delegatecall() might introduce. Ensuring that delegatecall() is only used with contracts declared as libraries is a solid practice.
III-E4 Example
Listing 11 demonstrates how an NFT fractionalization contract might be exposed to a code injection vulnerability through unsafe use of delegatecall. In this contract, executeDistributionLogic can execute arbitrary logic via delegatecall, based on the bytecode provided in data. If logicContract points to a malicious contract, it could lead to unintended alterations in the contract’s state, including token ownership manipulation.
To mitigate the risk of code injection via delegatecall, the contract should strictly control the update of the logicContract address and ensure that only verified, safe operations are executable. Also, we introduce a whitelist of approved logic contract addresses and require that any updates to logicContract come from this whitelist. Listing 12 illustrates this changes.
III-F Randomness Using Block Information
III-F1 Description
In blockchain-based applications, certain functionalities might require randomness—for instance, distributing a rare NFT fractionally among participants or deciding the winner of a unique NFT in a lottery, block information, such as block hash and block timestamp, can be used to achieve this. This information, however, may be anticipated and slightly modified by miners. When the block timestamp is used as the trigger condition to execute the transaction, it creates a vulnerable situation; indeed, dishonest miners can exploit the value of the block timestamp in an unethical manner. This vulnerability was exploited in GovernMental, a Ponzi scheme game [24]. The player who joined the round last and stayed for at least a minute was compensated according to the game rules. A miner who is also a player, might change the timestamp to make it look like he/she was the last to join for more than a minute and therefore collect the reward.
III-F2 Implications on NFT Fractionalization
Randomness Using Block Information could impact Fractionalization solutions. For instance, if randomness derived from block attributes decides the allocation of rare NFT fractions or the winners of rewards, it could lead to outcomes unfairly attribute in favor of those with the ability to influence block information. Similarly, auctions for selling fractionalized NFTs could be manipulated, allowing miners or others with insider advantages to affect the auction’s outcome by adjusting the voting period for example.
III-F3 Protection Measures
To address the vulnerabilities associated with using block information for generating randomness in smart contracts, adopting more reliable sources of randomness is essential. One solution is employing oracles [25] or other external sources that provide verified randomness is a robust solution.
Alternatively, cryptographic commitment schemes [26], exemplified by solutions like RANDOA [27] , represent another effective approach. These schemes involve participants committing to their inputs in a concealed manner, which are later revealed collectively to generate a random outcome. This method ensures that no individual can influence the result based on other participants’ commitments, fostering fairness and security.
III-F4 Example
Listing 13 shows an example of an NFT fractionalization platform that randomly assigns fractional shares of an NFT to participants based on block hash as a source of randomness. This contract is vulnerable because miners, or participants with mining capabilities, can potentially manipulate the block hash to influence the distribution outcome.
To mitigate this vulnerability, the platform can utilize an external oracle to provide a source of verified randomness. This approach ensures that the randomness used for fraction distribution is not predictable or manipulable by miners. Contract illustrated in Listing 14 inherits from VRFConsumerBase, provided by Chainlink [28], to securely request and receive verified random numbers. It uses the Chainlink VRF (Verifiable Random Function) to ensure that the randomness determining the distribution of NFT fractions cannot be influenced by participants, including miners.
III-G Other Vulnerabilities & Summary
Table II present the papers that include each vulnerability. We notice that the most mentioned vulnerability is reeantancy, whereas code injection is the least discussed.
Reentrancy | Front Running | Arithmetic | Mishandled exceptions | Code Injection | Randomness | |
[29] | X | |||||
[30] | X | X | X | |||
[31] | X | X | X | X | ||
[32] | X | X | X | |||
[33] | X | |||||
[34] | X | X | X | X | ||
[24] | X | X | X | X | X | |
[35] | X | X | X | X | ||
[36] | X | X | X | X | ||
[37] | X | X | X | X | ||
[5] | X | X | X | X | ||
[38] | X | |||||
[6] | X | |||||
[18] | X | X | X | |||
[39] | X | X | X | |||
[40] | X | X | ||||
[41] | X | |||||
[42] | X | |||||
[43] | X |
The Smart Contract Weakness Classification Registry (SWC) [44] is a community database of known smart contract vulnerabilities; it is often up to date. For each issue, it includes a description, code samples, and protection measures. Currently, the registry contains 36 vulnerabilities. One has to check this registry, often, to stay up to speed on the newest threats.
In addition, Rusinek et al. [45] proposed the Smart Contract Security Verification Standard (SCSVS). SCSVS is a 14-point checklist designed to standardize smart contract security for programmers, designers, security auditors, and vendors. By offering recommendations at every level of the smart contract development cycle, SCSVS helps avoid the most known security concerns and vulnerabilities.
IV Detection methods
In this section, we present the most common approaches for detecting smart contract vulnerabilities. They include static analysis, dynamic analysis, and formal verification. We will briefly describe each of these approaches and identify their benefits and limitations.
IV-A Static Analysis
Static code analysis is a technique of debugging that involves reviewing source code prior to running it; this technique is also known as white-box testing [46]. It is accomplished by comparing a set of code to a set of predefined coding rules. In fact, there are a number of techniques to examine static source code; they can be incorporated into a single solution. Compiler technologies are frequently used to develop these techniques such as Taint Analysis and Data Flow Analysis [47]. In this Section we present some static analysis techniques, namely Control Flow Graph, Taint Analysis, and Symbolic Analysis.
IV-A1 Control Flow Graph
The Control Flow Graph (CFG) is a directed graph that describes the flow of control. The flow of execution of source code A graph node denotes a basic block with no jumps; directed edges represent jumps from one block to another. If a node only has an exit edge, it is referred to as an entry. And if it has only an entry edge, it is referred to as an exit block. In Figure 15, Block 1448 is an entry block, and blocks 1451 and 1452 are exit blocks.
Figure 16 shows the common process to create CFG in the context of Ethereum. The first step is Parsing bytecode. Typically, the algorithm will extract the compiler version from the metadata; then, it will parse the remaining bytecode (without the metadata) into opcodes [48]. The parsing stage is straightforward because each of the bytecode’s two characters represent one opcode (for example 0x04 represents DIV). The complete list of opcodes can be found in the Ethereum yellow paper [49]. The second step is Identification of basic blocks. A basic block is a set of opcodes that run in succession between a jump target and a jump instruction, with no other instructions interrupting the control flow. To determine the basic blocks, we need to identify the opcodes that alter the execution of the control flow; some examples are presented in Table III. The JUMPDEST instruction starts a new basic block, whereas the other opcodes terminate blocks. Each block is identified by an offset that represents the position of its first opcode in the bytecode. The last step will be computing the edge by checking the destination offset; this step is not trivial as the destination is not an opcode parameter. In the literature, there are two types of jumps: (a) A pushed jump is a JUMP followed by a PUSH opcode, allowing the target offset to be determined simply by glancing at the data in the previous PUSH opcode; and (b) Orphan jump is not followed by a PUSH; indeed, its destination is not computed directly [50]. The computation of orphan jumps depends on each implementation and it is critical for the completeness of the graph.
Operation | Opcode | Description |
Flow | JUMP | Alters the program counter |
Alters the program counter | ||
Operations | JUMPI | Conditionally alters the program counter |
JUMPDEST | Marks a valid destination for jumps | |
System | STOP | Halts execution |
operations | REVERT | Halts execution reverting state changes |
but returning data and remaining gas | ||
RETURN | Halts execution returning output data | |
INVALID | Designated invalid instruction | |
SELFDESTRUCT | Halts execution and registers account | |
for later deletion |
IV-A2 Taint Analysis
Taint Analysis aims to detect input variables where data comes from an untrusted source that could be controlled by an attacker (e.g. environmental data or function parameters). These data inputs are tainted by the taint analysis tool and then traced back to potentially sensitive functions, such as security checks or storage access, often know as sinks. If the tainted variable flows into a sink, it is marked as a vulnerability [41].
In the context of Ethereum smart contracts, this technique is usually used after generating CFG. It exploits opcode that an attacker may employ to insert data into the control flow; it can be divided into block information, environmental information, and flow information. Table IV shows an example of opcode that allows data insertion can be found in .
Operation | Opcode | Description |
Block | GASLIMIT | Gets the current block gas limit |
information | TIMESTAMP | Gets the current block timestamp |
Environment | CALLER | Gets caller address |
information | CALLDATALOAD | Gets input data of current |
environment | ||
CALLDATACOPY | Copies input data in current | |
environment to memory | ||
Flow | SLOAD | Loads word from storage |
Operations | MLOAD | Loads word from memory |
An example of taint analysis is to detect arithmetic bugs. The first step is to define the possible entrusted inputs (source) such as CALLDATALOAD and SLOAD. The affected memory, storage or stack location is tagged. The next step is to check whether the flow of tagged input contains arithmetic operations such as ADD, MUL and SUB with no catch mechanism (sink). For example, a catch mechanism is reverted in case of DIV with 0. If the tagged input falls into a sink, it will be marked as a possible vulnerability.
IV-A3 Symbolic Analysis
Symbolic execution is a way of abstracting the execution of a program such that it can span numerous pathways through the code. The program is run with symbols as inputs, and the outputs are expressed in the form of the symbolic inputs. Each symbolic path has a condition; it is a formula created by collecting constraints that must be satisfied by those inputs in order for the execution to continue on that path. If the condition is unsatisfiable, the path is infeasible; otherwise, the path is feasible [51]. Symbolic analysis is usually paired with CFGs. In order to find exploits, most tools employ the Satisfiability Modulo Theories (SMT) [52] solver to check whether the symbolic output is feasible or provide a counterexample if it is not.
IV-B Dynamic Analysis
The term ”dynamic analysis” refers to the observation of code while it is being executed in its original context. It is also known as black-box testing because it is performed without access to the source code. It works in the same way as an attacker who feeds malicious code or unpredictable input to the appropriate functions of a program to look for vulnerabilities. The most common technique is Fuzzing.
IV-B1 Fuzzing
Fuzzy testing, often known as fuzzing, is a kind of automated software testing that involves injecting incorrect, malformed, or unpredictable inputs into a system; the objective is to uncover software flaws and vulnerabilities. A fuzzing tool injects these inputs into the program and then watches for problems such as crashes or data leaks; it also can perform static analysis on execution traces. In the case of smart contracts, Wang et al. [53] deploy the smart contract on a test network; then, they monitor the balance of the smart contract to identify basic misappropriation. This technique removes the necessity for particular patterns to identify a vulnerability.
IV-C Formal Verification
Formal verification is a technique that automates bug detection of a hardware or software system by comparing a formal model of the system to formal requirements or behaviour specifications. Through mathematical analysis, formal verification can provide a high level of confidence.
In order to conduct a formal verification, first we need to provide a specification. A program specification is an unambiguous definition the purpose of the program as well as the scenarios that are allowed or not to be executed. In order to verify a property, tools usually build deduction trees where the root presents a Hoare triplet [54]. A Hoare triplet consists of: (a) Pre condition: It is the initial state; (b) Instructions: A series of transactions; and (c) Post condition: It is the property to verify. In order to build deduction trees theorem proving algorithms are usually used such as Coq [55] and Isabelle/HOL [54].
There are a few contributions that propose languages to write formal specifications. Permenev et al. [56] proposed a specification language, called VerX, where the syntax is similar to Solidity; it supports temporal logic. The Ethereum community also proposed a language named Act [57] which is a specification for Ethereum Virtual Machine (EVM) programs.
IV-D Comparison
The goal of using detection methods such as static, dynamic, and formal verification is to ensure that smart contracts are correctly executed in all states; this is to ensure that no vulnerabilities are produced to ensure that specifications are respected. Table V shows the papers that discuss each methodology.
Static code analysis tools can potentially produce false negatives results; vulnerabilities occur but are not reported. This could happen because the analysis tool used doesn’t allow for a large number of test scenarios; it doesn’t go into detail when looking at different states because it usually gives you a timeout. Various vulnerabilities may be false negatives in static analysis; however, they can be detected correctly utilizing dynamic analysis tools. Thus, it is strongly recommended to combine both types of analysis.
Dynamic analysis generates transactions with random inputs. This means that it will observe random states depending on inputs. The challenge for these techniques is to generate enough transactions and to carefully select inputs to achieve maximum coverage.
Formal verification can provide full coverage for the specification, under consideration, because it employs mathematical analysis. This guarantees the satisfaction of a property if verified; however, in general, we cannot cover everything because the verification can become easily intractable. Thus, false negatives can still be present. In addition, formal verification requires a skilled programmer who can express a smart contract as a mathematical high-level specification while accounting for a specific low-level virtual machine. This operation takes a long time and requires a lot of resources. Audit firms usually make use of formal verification [58].
V Vulnerability Detection Tools
In this section, we will cover the most popular tools for detecting smart contract vulnerabilities. To investigate available tools, we first looked through academic literature and review articles (see Table VI). Then we covered five of the most commonly used tools as indicated by developer forums as well as the number of forks and update frequency on GitHub.
# | Tool | Detection Technique | Last update |
1 | Slither [6] [67] | Static Analysis | May 2023 |
2 | MythX [68] | static, dynamic & Symbolic Execution | Not Public |
3 | Mythril [7] [69] | Symbolic Execution | May 2023 |
4 | Echidna [9] [70] | Fuzzer | Apr 2023 |
5 | Manticore [8] [71] | Symbolic Execution | June 2022 |
6 | Securify [39] [72] | Static Analysis | Sep 2021 |
7 | KEVM [60] [73] | Static Analysis | Apr 2022 |
8 | Smartcheck [18] [74] | Static Analysis | Dec 2019 deprecated |
9 | MadMax [42] [75] | static analysis | Jun 2021 |
10 | Vertigo [76] [77] | Mutation Testing | Feb 2021 |
11 | EtherSolve [38] [78] | CFG Extraction | Nov 2021 |
12 | Octopus [79] | Symbolic Execution | Nov 2020 |
13 | Oyente [5] [80] | Symbolic Execution | Nov 2020 deprecated |
14 | ERC20 Verifier [81] | Verify ERC20 Compatibility | Nov 2019 |
15 | Solgraph [82] | CFG Extraction | Jan 2019 |
16 | Osiris [41] [83] | Symbolic Execution | Sep 2018 |
V-A Oyente
V-A1 Description
Luu et al. propose Oyente [5] the first symbolic execution tool for Ethereum smart contracts. Oyente was the basis of several tools developed later [62] [41]. It has four components: (a) CFGBuilder: It constructs CFG of the smart contract; (b) Explorer: It takes as input the Ethereum state. It has a loop that runs a state then executes an instruction on the output of that state. The loop continues until no state remains or a timeout is reached. The output is a symbolic trace. Explorer determines the infeasible trace by querying the Z3 SMT solver [84]; (c) CoreAnalysis: It targets predefined vulnerabilities by looking for patterns on the symbolic trace; (d) Validator: It queries Z3 solver with traces flagged as vulnerable to reduce false positives cases.
Oyente detects seven types of vulnerabilities, namely Re-entrancy, Integer overflow/underflow, Transaction order dependence, Timestamp dependence, Callstack Depth, EVM Code Coverage, and Parity Multisig bugs.
V-A2 Execution Example
For evaluating Oyente’s capabilities in the context of smart contracts dealing with token sales and potentially fractionalized assets, we utilize the TokenSaleChallenge.sol smart contract, sourced from the Capture the Ether challenge series [85]. The contract embodies a token sale mechanism where tokens are bought and sold at a fixed price, presenting an analogous scenario to NFT fractionalization where individual tokens could represent shares of a fractionalized NFT.
The primary aim is to assess Oyente’s effectiveness in detecting vulnerabilities that could also impact contracts handling fractionalized NFTs. The results, depicted in Figure 17, highlight an Integer Overflow vulnerability, along with the execution trace where this issue is detected.
V-A3 Pros and Cons
Oyente’s deployment via a Docker image significantly simplifies its setup process, making it accessible for users with varying levels of expertise in blockchain technology. This ease of use facilitates quick integration into development workflows, allowing for the immediate analysis of Ethereum smart contracts. Despite its user-friendly setup, Oyente’s compatibility is somewhat limited; it supports Ethereum compiler versions only up to 0.4.17. This restriction may hinder its applicability to smart contracts compiled with newer versions of the Solidity compiler, potentially limiting its utility in analyzing the latest smart contract developments.
The output provided by Oyente is both clear and concise, enabling developers and auditors to easily interpret vulnerability reports. Moreover, Oyente has played a pivotal role in the evolution of smart contract analysis tools. It has laid the groundwork for subsequent innovations, including Maian [86] and Osiris [41]. These tools have built upon Oyente’s foundational techniques, such as the construction of Control Flow Graphs (CFGs), to offer enhanced accuracy in detecting vulnerabilities and orphan paths within smart contracts.
V-B Slither
V-B1 Description
In their work, Feist et al. [6] unveiled a static analysis instrument known as Slither. This tool is aimed at spotting weaknesses, enhancing code efficiency, and augmenting the understanding of code. The procedure it employs is in multiple stages: (a) Data Retrieval: Initially, it creates the Abstract Syntax Tree of the contract’s source code using the Solidity compiler, yielding some information like the Control Flow Graph and contract inheritance; (b) Conversion into SlitherIR: The source code of the smart contract is then converted into an internal representation language referred to as SlithIR; and (c) Examination of the Code: This phase involves determining the variables being read and written as well as their types. Furthermore, it detects unsecured functions where potentially harmful addresses can execute high-level operations. During this Examination of the Code stage, it also calculates data dependency and marks variables that rely on user input as tainted.
Slither is proficient in uncovering 70 bug varieties, which encompass self-destructive smart contracts, code injections via delegatecall, frozen ether, and Reentrancy vulnerabilities [87].
V-B2 Execution Example
We analyzed the FractionalNFTMarket smart contract in Listing 18 using Slither to identify potential security vulnerabilities. The results of the analysis are presented in Figure 19. Slither has successfully detected a reentrancy vulnerability within the sellShares function of the contract, which could allow an attacker to exploit the contract by recursively calling the function to drain its funds. In addition to detecting this critical vulnerability, Slither offered recommendations for code improvements and provided a comprehensive set of information, including a human-readable summary, an inheritance graph, and the Control Flow Graph (CFG) of each function.
Slither also provides a list of useful information including a human readable summary of the scanned contract, inheritance graph, and CFG of each function; the complete list can be found in the github repository [67].
V-B3 Pros and Cons
Slither, an open-source static analysis tool, is recognized for its efficiency in auditing smart contracts. It is user-friendly, offering straightforward installation options via Docker or Python package managers. With the capability to detect approximately 70 types of vulnerabilities, Slither facilitates improved code understanding through visual aids like contract graphs. Moreover, it examines smart contract compliance with established Ethereum Request for Comments (ERC) standards, including ERC-20 and ERC-777 [88].
One limitation of Slither is its tendency to report false positives, which can complicate the auditing process by requiring additional validation to go through the findings. For a more comprehensive analysis, Crytic [89], a premium service, can be utilized. Crytic extends Slither’s capabilities by identifying an additional 50 types of faults that Slither might not catch. Furthermore, Crytic offers integration with GitHub, allowing for automated testing on pull requests, thereby enhancing the development workflow and ensuring continuous contract integrity.
V-C Mythril
V-C1 Description
Slithe Durieux et al. [7] developed a tool known as Mythril that is capable of identifying potential weaknesses in smart contracts by utilizing symbolic execution. Mythril, a Python-based command-line tool, allows for interactive inspection of smart contracts. It produces a Control Flow Graph (CFG) and performs symbolic execution of EVM bytecode to restrict the search area. The tool can generate concrete values to exploit any detected vulnerabilities. To evaluate the feasibility of different paths, Mythril makes use of the Z3 SMT solver [84].
Mythril has the ability to discover fourteen distinct types of vulnerabilities including Delegate Call To Untrusted Contract, Dependence on Predictable Variables, Deprecated Opcodes, Ether Thief, Exceptions, External Calls, Integer Over/Underflow, Multiple Sends, Self-Destruction, State Change External Calls, Unchecked Return Values, User Supplied Assertions, and Arbitrary Storage Write and Arbitrary Jump [90].
V-C2 Execution Example
The FractionalNFTMarket.sol smart contract, presented in Listing 18, has been analyzed using Mythril to uncover potential security issues. The resulting output is depicted in Figure 20. Mythril has identified a reentrancy vulnerability within the contract’s ‘sellShares‘ function. The output provides detailed information, including the state of the contract when the vulnerability can be triggered and the specific transaction sequence that leads to the issue. Unlike Slither, Mythril offers not only the identification of vulnerabilities but also the concrete inputs that cause them, enhancing the developer’s understanding of the contract’s weaknesses. Nevertheless, Mythril’s output does not extend to include recommendations for remediation of the detected issues.
V-C3 Pros and Cons
Mythril stands out for its accessibility, offering straightforward installation options via Docker or Python package manager, which streamlines the setup process. It allows for granular analysis, with the flexibility to test individual functions through custom Python scripts. This targeted testing can be particularly advantageous for developers seeking to debug specific aspects of their smart contract code.
However, Mythril’s thorough approach to smart contract analysis can be computationally intensive, often resulting in longer analysis times compared to tools like Slither. The resource-heavy nature of Mythril’s symbolic execution process means that it may not be the most efficient choice for rapid iteration during development. For those seeking a more advanced feature set and cloud-based capabilities, Mythril’s commercial counterpart, MythX [68], offers enhanced analysis power and can be seamlessly integrated into continuous integration/continuous deployment (CI/CD) pipelines [91].
V-D Manticore
V-D1 Description
Mossberg et al. [8] implemented a tool named Manticore to detect vulnerabilities in smart contracts. The primary aim of Manticore is to track inputs that kill a program, record instruction-level implementation, and provide Python API access to its analysis engine. To locate distinct computation paths in EVM bytecode, Manticore uses symbolic execution. It discovers inputs that will stimulate these computation paths with the aid of the SMT solver Z3 [84]. It keeps track of the execution traces for each execution. Manticore converts Solidity code to bytecode for evaluation; then, it examines the traces for vulnerabilities, such as reentrancy and reachable self destruct operations, reporting them in the context of the source code. This tool is developed by TrailOfBits [92].
Mythril detects ten types of vulnerabilities namely, delegatecall, overflow, reentrancy, Use of potentially unsafe/manipulable instructions, Reachable external, Reachable selfdestruct instructions, Uninitialized memory usage, Uninitialized storage usage, Enable INVALID instruction detection, and Unused internal transaction return values [93].
V-D2 Execution Example
For our examination of Manticore, we used the TokenSaleChallenge.sol [85] same one used to test Oyente. This contract represents a simplified model for trading tokens, which can be seen as stand-ins for fractional tokens of an NFT.
The tool is extremely slow as it takes more than one hour to perform the test; it takes less than a minute for the other tools using the same smart contract and machine. The output of the analysis is a folder containing a summary report as well as traces and analysis for each test case. Manticore generated 51 test-cases, but it failed to detect the smart contact’s reentrancy vulnerability, as shown in Figure 21.
V-D3 Pros and Cons
Manticore offers a diverse set of command-line tools along with scriptable Python APIs, which is suitable to a variety of use cases and allow for in-depth analysis procedures. Manticore’s cover a broad spectrum of known vulnerabilities, making it a valuable asset for developers seeking to secure their smart contracts.
However, the detailed nature of Manticore’s analysis comes with a trade-off in terms of performance. The tool can take significantly longer to complete its analysis compared to other tools, sometimes leading to timeouts. This is often due to its comprehensive symbolic execution approach, which, while powerful, is also resource-intensive. Additionally, Manticore is known to require a considerable amount of memory, which could be a limiting factor for users with constrained hardware resources.
V-E Echidna
V-E1 Description
Echidna is a specialised property-based fuzzing tool for smart contracts implemented by Grieco at al. [9] that takes inspiration from QuickCheck [94]. Echidna seeks to violate user-defined invariants that represent potential contract errors rather than just finding crashes [95]. These invariants are writtern by the tester and it is called echidna property: A special Solidity function with no arguments that returns ”true” on success and has a name that begins with ”echidna”. Any transactions that result in these properties returning false or error are reported by Echidna, essentially disclosing contract bugs [96]. Echidna employs a variety of strategies, including iterative feedback, structural constraints, and known input variation, to produce test inputs.
V-E2 Execution Example
In order to test Echidna, we have adapted the tokensalechallenge.sol smart contract [85]. We created a derived contract TestTokenSaleChallenge that extends the original TestTokenSaleChallenge contract shown in Listing 22.The test properties in the TestTokenSaleChallenge are designed to check for arithmetic errors that can occur in the logic used for NFT fractionalization.
The echidna_test_overflow function aims to validate that token balances do not exceed the maximum uint256 value, a critical check to prevent overflow in token allocations, which could mirror similar concerns in the distribution of fractional NFT shares. Conversely, echidna_test_underflow ensures that the balance never drops below zero.
By defining these test functions, Echidna will attempt to generate test inputs that trigger these conditions to ensure the contract behaves as expected.
Echidna generates its results directly to the console, as figure 23 shows, distinguishing between properties that were successfully validated and those that failed. For each property that didn’t meet the validation criteria, Echidna offers a counterexample, which is essentially a step-by-step breakdown of how the property failed under specific conditions. In our example, the underflow test did not pass, even without initiating a transaction. This was because the msg.sender was left empty, indicating that the system can encounter failures in scenarios where sender information is not provided. However Echidna didn’t detect an overflow issue, possibly due to a short timeout period, preventing it from producing a sequence revealing this vulnerability as we did not change the default configuration.
V-E3 Pros and Cons
One of Echidna’s notable strengths lies in its straightforward setup process, offering users multiple installation options, including direct build with Stack, Docker, or Homebrew. This flexibility facilitates accessibility across different platforms and development environments.
However, users are required to explicitly define the properties they wish to test. This feature allows for customized and focused testing scenarios, ensuring that the analysis is directly relevant to the contract’s intended behaviors and security assumptions. Despite the advantage of tailored testing, this approach adds a layer of complexity. Users must possess a thorough understanding of their contract’s logic and potential vulnerabilities to effectively define meaningful test properties. This necessity for manual property definition demands a higher level of engagement and expertise from the user, compared to tools that automatically infer test cases or vulnerabilities.
V-F Summary
Table VII presents a comparative analysis of Oyente, Slither, Mythril, Manticore and Echidna. The comparison is based on the number of detected vulnerabilities, used methodology, an academic or company tool, the code level, required solidity version, and availability of documentation.
Oyente | Slither | Mythril | Manticore | Echidna | |
Number of detected vul | 4 | 70 | 14 | 10 | - |
Methodology | Symbolic | Symbolic | Static | Symbolic | Fuzzing |
Code Level | EVM | Solidity | Solidity | EVM | Solidity |
Restrictions (Solidity) | 0.4 | 0.4 | 0.4 | 0.4 | not specified |
Documentation | Low | Medium | High | High | Medium |
In conclusion, there are a variety of tools that can be used to test smart contracts and identify a wide range of vulnerabilities. Static analysis tools are still the most popular because they are simple to use compared to (a) fuzzing which necessitates a lot of resources to set up the test environment; and (b) formal formal verification which necessitates expertise in contract specification writing.
VI Guidelines for secure smart contracts
Writing a secure and bug-free smart contract is crucial as a small bug can lead to the loss of millions of dollars. In this section, we provide guidelines on how to write a secure smart contract [97]. These steps are summarized in Figure 24
The first step to test a smart contract is to visualize its control flow right in order to have a global view and better understanding of the interactions between its different components. Slither is an interesting tool to use during this step since it has a variety of printers that can be used to visualise call graphs, contract inheritance relationships, modifiers called by each function, etc. Solidity Visual Developer is another interesting tool that could be integrated into Visual Studio Code [98]. This extension helps with semantic highlighting. It also generates a detailed class outline. This step is important to spot critical functions that must be tested in the next stage.
Afterwards, the smart contract should be subjected to an automatic analysis. It is best to start with common bug detection using Mythril and Slither, we recommend combining tools for better results. To target a certain critical function, it is best to use Manticore tool as it allows the implementation of specific use cases. It is also crucial to employ dynamic analysis tools to decrease the number of false positive generated by static tools and to cover additional states. Echidna is an interesting ethereum fuzzing tool to use [9]. It is recommended to look for unique features in smart contracts. For instance, if the smart contract is an ERC token, it will be important to use Slither’s ERC Conformance functionality to make sure that the smart contract conforms to the ERC standards put in place.
Because automated tools are unaware of the context for which the contract was produced, running more specific testing will increase the detection accuracy. To achieve this, Echidna and Manticore [99] allow for the definition of security properties in solidity. Indeed, these tools are useful for checking arithmetic operations, external interactions, and standardisation. Whereas Slither allows the definition of properties with Slither Python API’s, it allows checking for inheritance, variable dependencies, and access restrictions.
The third step will be formal verification; it is crucial since it ensures that the contract requirements are legitimate. This step, however, is challenging as it requires special expertise (e.g., writing the formal specification using a specification language). Finally, the output of the preceding steps must be manually verified and corrective actions based on patterns and mitigation techniques must be applied. After fixing the bugs, one has to redo the previous steps to increase the probability of a bug free contract.
VII Related Work
In the last few years, several of surveys covering smart contract vulnerabilities have been published. The majority of these surveys cover one or two of the following areas: Security vulnerabilities of smart contracts, analysis of existing tools and methodologies to detect vulnerabilities of smart contracts, formal specification and verification, and common design patterns and practices of smart contracts. Kushwaha et al.[100] conducted a systematic review of Ethereum smart contract analysis tools, classifying them into static and dynamic analysis categories. They scrutinize 86 tools, discussing their methodologies, such as taint analysis, symbolic execution, and fuzzing. However, it lacks of consideration for the developer community’s preferences and practical scenarios where each analysis tool would be most beneficial. Kushwaha et al. [21] presented a comprehensive systematic review of security vulnerabilities in Ethereum blockchain smart contracts. The paper categorizes vulnerabilities into three main root causes and seventeen sub-causes, providing in-depth insights into twenty-four specific vulnerabilities along with their prevention methods, detection, and analysis tools. Despite providing a comparative analysis of various detection tools, the paper falls short in offering in-depth feedback on the effectiveness and limitations of these tools. Di Angelo et al. [101] provided a state-of-the-art review of analysis tools of Ethereum smart contracts. The study is based on real execution of the tools. It classifies the tools based on availability, maturity level, purpose, and analysis method. However, the survey [101] become obsolete (e.g., no longer maintained). Harz et al. [102] presented languages, paradigms, and a verification approach for smart contracts. They did not; however, go into depth about the verification tools or known vulnerabilities. Li et al. [103] discussed blockchain safety problems and proposed enhancements, however, the survey [103] is general because, for example, it included different types of blockchain, such as Ethereum and bitcoin. Saad et al. [104] investigated several attacks on various blockchain platforms, as well as protection mechanisms. They briefly covered mitigation schemes, however, they did not cover tools to detect vulnerabilities. Atzei et al. [24] presented major vulnerabilities and attacks related to Ethereum smart contracts, however, they did not cover mitigation schemes. Chen et al. [105] presented 40 types of Ethereum vulnerabilities, along with their causes and 29 attacks; however, they did not cover tools to detect vulnerabilities. Zhu et al. [106] presented 11 smart contract vulnerabilities in depth, along with various defences against well-known attacks. They did cover schemes to detect those vulnerabilities. Durieux et al. [107] conducted an empirical review of nine automated analysis tools on 47,518 contracts, They developed a framework to analyze these tools. The analysis makes use of two data sets of smart contracts with tagged vulnerabilities. However, the study did not cover vulnerabilities of smart contracts and mitigation schemes.
We conclude that existing surveys focus only on one or two areas of smart contract vulnerabilities. Furthermore, none of the studies discuss the tools based on their output format.
VIII Conclusion
In this study, we presented detailed common vulnerabilities in Ethereum smart contracts, as well as mitigation solutions based on patterns and standards. We have covered the most popular detection methodologies, such as static analysis, dynamic analysis, and formal verification. We also did discuss the benefits and drawbacks of each methodology. In addition, we provided community-recommended vulnerability detection tools, along with an execution sample and detailed feedback for each. Based on our investigation of existing methodologies and tools, we proposed a guideline with the tool(s) to be used for each step in auditing smart contracts.
We did observe that most of the tools detect vulnerabilities but do not provide refactoring recommendations. In future work, we aim to implement a refactoring module that proposes improvements based on the detected bugs and error traces. We also aim in study attacks related to NFTs such as wash trading which is a type of market manipulation where attackers repeatedly buy and sell the same NFT to increase the price and trading volume.